OpenStudy (anonymous):

I GIVE MEDALS. A gargoyle statue falls off the side of a building from a height of 75 m. (Assume no air resistance.) How long does it take the statue to reach the ground? I think i use the equation d=1/2(a)(t2)+v1(t) but im not sure where to plug in all the numbers to solve.

OpenStudy (anonymous):

You have the right formula, you just need to plug in d = 75 m a = 9.8 ms^-2 v1 = 0

OpenStudy (anonymous):

how would the equation look rearranged so that it's t=? @broken_symmetry

OpenStudy (anonymous):

Well let's start with \[d = \frac12 a t^2\] Multiply both sides by 2/a \[\frac{2d}a = t^2\] take the square root of both sides \[\sqrt\frac{2d}a = t\]

OpenStudy (anonymous):

so the final answer would be 3.9? @broken_symmetry

OpenStudy (anonymous):

Yeah http://www.wolframalpha.com/input/?i=75+%3D+4.9*x^2

OpenStudy (anonymous):

how would i figure out how fast is the statue moving when it reaches the ground? @broken_symmetry

OpenStudy (anonymous):

Well, you know the acceleration (9.8 ms^-2) and you just worked out the time. You know it was stationary to begin with. So...

OpenStudy (anonymous):

is it the same equation? @broken_symmetry

OpenStudy (anonymous):

No, that equation was to get distance. The equation for velocity from acceleration and time is much simpler.

OpenStudy (anonymous):

is it just displacement divided by time? @broken_symmetry

OpenStudy (anonymous):

That gets you the average velocity, but you want the final velocity. Acceleration is velocity divided by time. There's a clue in the units \[\text{m s}^{-2}\] So if something has been accelerating at 9.8 ms^-2 for 3.9 s, how fast is it now going?

OpenStudy (anonymous):

it would just be 38.22 correct? @broken_symmetry

OpenStudy (anonymous):

Yep :)

OpenStudy (anonymous):

Just a word of advice: I found it really really useful to memorise the SUVAT equations: \[v = u+at\] \[s=ut+\frac12at^2 = vt-\frac12at^2\] \[v^2=u^2+2as\] \[s=\frac{u+v}{t}\] s is distance, u is initial velocity, v is final velocity, a is acceleration and t is time. You can choose which one to use by which one you know all the variables except the one you want to work out.

OpenStudy (anonymous):

thanks so much! @broken_symmetry